# Deep Legendre Transform

### v1

@@ B.3.2 Benchmark setup
Table B8: Benchmark results comparing implicit DLT against direct learning with known duals
Function $d$ Model $L^{2}$ Error Time (s) Ratio
Impl. Dir. Impl. Dir. $\rho$ $\sigma$
Quadratic 2 MLP 1.70e-05 1.30e-05 69.5 78.5 1.42 0.55
MLP_ICNN 4.44e-01 5.03e-01 66.1 62.7 1.54 2.22
ResNet 2.24e-05 1.03e-05 75.5 94.3 2.51 2.56
ICNN 8.26e-04 8.79e-04 91.0 91.2 1.04 0.62
5 MLP 4.19e-04 4.13e-04 81.0 81.3 1.07 0.40
MLP_ICNN 3.23e-01 3.18e-01 82.8 83.4 1.02 0.08
ResNet 2.03e-04 2.05e-04 93.3 93.3 1.02 0.25
ICNN 1.82e-03 1.80e-03 89.9 89.8 1.02 0.22
10 MLP 5.68e-03 5.64e-03 81.1 77.9 1.02 0.19
MLP_ICNN 4.74e-01 4.81e-01 83.6 83.6 0.99 0.02
ResNet 2.03e-03 2.06e-03 92.6 92.5 1.00 0.17
ICNN 6.45e-03 6.18e-03 89.7 89.6 1.06 0.15
Neg. Log 2 MLP 7.32e-05 6.08e-05 89.7 94.2 1.20 0.62
MLP_ICNN 5.33e-01 5.20e-01 77.1 83.1 1.06 0.28
ResNet 1.42e-05 1.37e-05 99.3 103.8 1.23 0.68
ICNN 2.76e-04 3.19e-04 96.6 101.1 0.93 0.25
5 MLP 1.68e-03 1.47e-03 90.6 95.2 1.17 0.32
MLP_ICNN 1.24e+$0$ 1.22e+$0$ 83.7 88.0 1.04 0.25
ResNet 6.24e-04 6.74e-04 100.2 104.8 0.95 0.24
ICNN 1.64e-03 1.73e-03 96.9 101.5 1.06 0.56
10 MLP 8.36e-03 8.31e-03 89.8 95.1 1.04 0.18
MLP_ICNN 3.06e+$0$ 3.04e+$0$ 78.6 79.9 1.01 0.05
ResNet 4.85e-03 4.44e-03 100.1 104.7 1.10 0.18
ICNN 2.56e+$0$ 2.12e+$0$ 82.3 81.6 201.30 403.37
Neg. Entropy 2 MLP 6.09e-05 4.63e-05 89.8 96.1 1.39 0.33
MLP_ICNN 1.19e-02 1.18e-02 90.4 99.2 1.22 0.78
ResNet 5.40e-05 5.00e-05 95.8 102.3 1.27 0.71
ICNN 2.22e-04 2.02e-04 96.2 105.1 1.15 0.26
5 MLP 3.41e-03 3.29e-03 88.2 96.6 1.07 0.24
MLP_ICNN 3.55e-02 3.63e-02 90.8 98.7 0.99 0.12
ResNet 1.63e-03 1.79e-03 99.8 108.4 0.95 0.23
ICNN 7.70e-04 6.91e-04 96.1 105.7 1.27 0.61
10 MLP 2.25e-02 2.03e-02 90.1 99.2 1.12 0.14
MLP_ICNN 7.85e-02 7.93e-02 90.7 99.4 0.99 0.08
ResNet 1.53e-02 1.46e-02 100.0 108.3 1.04 0.22
ICNN 2.06e-03 2.06e-03 97.0 105.4 1.01 0.21
@@ B.4 Experiments with non-separable functions
Pre-trained ICNN: We first train a 2-layer ICNN (128 hidden units per layer) for 50,000 iterations to approximate a target quadratic function, then freeze it and use it as the convex function $f$. This represents a realistic scenario where the target function itself is a learned model from data.
@@
(pre-trained to quadratic) ($\approx[-3,3]^{d}$ )
@@
Table B10: Performance of DLT with ResNet approximators on non-separable functions. Absolute RMSE, relative RMSE (normalized by function scale), and training times are reported as mean $\pm$ standard deviation over 5 independent trials.
@@ Results.
The results demonstrate that DLT successfully learns all four non-separable function classes, achieving low RMSE errors across different dimensions. The quadratic SPD function shows the best performance with errors decreasing as dimension increases (from $1.15\times 10^{-2}$ at $d=10$ to $2.01\times 10^{-3}$ at $d=50$), demonstrating DLT’s effectiveness in learning structured covariance patterns. The exponential-minus-linear function presents greater challenges due to its unbounded gradient domain, yet DLT maintains reasonable accuracy. The pre-trained ICNN case validates DLT’s ability to approximate learned neural network functions.
@@
B.5 Alternative NN approaches to computing Legendre transform
Our Deep Legendre Transform (DLT) approximates the convex conjugate $f^{*}$ on $D=\nabla f(C)$ by leveraging the exact Fenchel-Young identity at gradient-mapped points. The key insight is that for any $x\in C$ with $y=\nabla f(x)$, the identity $f^{*}(y)=\langle x,y\rangle-f(x)$ holds exactly.
@@ B.5 Alternative NN approaches to computing Legendre transform
The DLT approach provides several critical advantages over the Proxy method:
@@
    Efficient dampling: DLT uses the exact gradient mapping $\nabla f$ for sampling, which is computationally efficient and exact. The Proxy method requires every new sample to pass through the learned inverse network $h_{\vartheta}$, adding computational overhead and potential approximation errors at inference time.
@@
    Training overhead: While both methods may use an inverse mapping for sampling control, the Proxy method critically depends on the quality of this mapping for its training targets, requiring more careful and time-consuming pre-training.
Table B11: DLT with approximate-inverse sampling vs. proxy. $h_{\vartheta}$ RMSE measures pre-learned inverse quality; $t_{h}$: inverse pre-training time.
@@ B.6 Hamilton–Jacobi equation and time-parameterized inverse gradient sampling
$\begin{cases}\partial_{t}u+H(\nabla_{x}u(x,t))=0,&(x,t)\in Q,\\ u(x,0)=g(x),&x\in\mathbb{R}^{d},\end{cases}$ (B1)
@@ B.6.1 The Hopf formula and function definitions
The Hopf formula provides an analytical solution to the Hamilton–Jacobi equation:
$u(x,t)$ $=(g^{*}+tH)^{*}(x),$ (B2)
$=\inf_{p\in\mathbb{R}^{d}}\left\{g^{*}(p)+tH(p)-\langle x,p\rangle\right\}.$ (B3)
@@
These provide a more challenging test case with nonlinear behavior:
@@
Both the assumptions of the Hopf formula: they are coercive and everywhere well-defined.
@@ B.6.2 Time-parameterized inverse gradient sampling
The core computational challenge in implementing the Hopf formula lies in solving the optimization problem in (B3). Our Time-DLT approach requires sampling pairs $(x,p)$ where $p=\nabla u(x,t)$ represents the solution gradient. This sampling strategy depends on the function type and the time parameter.
Direct sampling for quadratic functions.
The quadratic case admits an explicit gradient relationship:
$p=\nabla u(x,t)=\frac{x}{1+t},$
allowing efficient direct sampling. Given any point $x$, we can immediately compute the corresponding gradient pp.
Inverse gradient network for exponential functions.
For the cases when the gradient relationship is implicit and complex (like computing $g^{*}+tH$ in the non-quadratic case), we train a time-parameterized neural network $\Psi(p,t):\mathbb{R}^{d}\times\mathbb{R}^{+}\to\mathbb{R}^{d}$ that learns the inverse mapping:
$\Psi(p,t)\approx x\quad\text{such that}\quad p=\nabla u(x,t).$
The network training employs a two-phase approach with complementary objectives:
  1. 1.
    Forward consistency: $\Psi(\nabla u(x,t),t)\approx x$
  2. 2.
    Inverse consistency: $\nabla u(\Psi(p,t),t)\approx p$
This unified network handles all time values simultaneously, learning the temporal evolution of the inverse gradient mapping.
B.6.3 Time-parameterized DLT implementation
Our Time-DLT method directly approximates the Hopf formula through a neural network $F_{\theta}(x,t)$ that learns:
$F_{\theta}(x,t)\approx u(x,t)=\inf_{p}\left\{\langle x,p\rangle-g^{*}(p)-tH(p)\right\}.$
@@ B.6.3 Time-parameterized DLT implementation
$\mathcal{L}=\mathbb{E}_{(x,p,t)}\left[\left|F_{\theta}(x,t)-\left(\langle x,p\rangle-g^{*}(p)-tH(p)\right)\right|^{2}\right].$
The sampling of $(x,p,t)$ triplets depends on the function type: analytical formulas for quadratic functions and the trained inverse network for exponential functions. This approach preserves the structure of the Hopf formula while ensuring computational tractability.
@@ B.6.4 Experimental framework
We evaluate both methods across increasing dimensions and function complexities.
@@ B.6.5 Implementation details
To evaluate accuracy, we compute reference solutions via numerical optimization of the Hopf formula (B3):
@@
  • Higher dimensions ($d>2$): PyTorch gradient-based optimization with Adam
@@ B.6.6 Evaluation metrics
    $\epsilon_{L^{2}}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}|u_{\theta}(x_{i},t)-u_{\text{ref}}(x_{i},t)|^{2}}$ (B4)
@@
    $\epsilon_{\text{PDE}}=\sqrt{\mathbb{E}_{(x,t)}\left[|\partial_{t}u_{\theta}+H(\nabla_{x}u_{\theta})|^{2}\right]}$ (B5)
@@
    $\epsilon_{\text{IC}}=\sqrt{\mathbb{E}_{x}\left[|u_{\theta}(x,0)-g(x)|^{2}\right]}$ (B6)
@@
Table B12: Comparison of methods for solving the Hamilton-Jacobi equation with Quadratic Initial Condition and Hamiltonian
@@
L2 Error PDE Res. Time (s) L2 Error PDE Res. Time (s) Mean $\sigma$
@@
Notes: L2 Error represents the mean squared difference between predicted solutions and the analytical solution $u(x,t)=\frac{\sum x_{i}^{2}}{2(1+t)}$. HJ Error measures the residual of the Hamilton-Jacobi equation $|\partial_{t}u+H(\nabla_{x}u)|$.
Table B13: Comparison of DGM and Time-DLT for solving the Hamilton–Jacobi equation with Quadratic Initial Condition and Exponential Hamiltonian
@@
L2 Error PDE Residual Time (s) L2 Error PDE Residual Time (s) DGM / DLT
@@
Notes: L2 Error quantifies the mean squared deviation from the ground truth solution. PDE Residual is the residual of the Hamilton–Jacobi equation $|\partial_{t}u+H(\nabla_{x}u)|$.
Figure B1: Comparison of DGM and Time-DLT across three different moments of time $t=0.1,0.5,1$, $d=2.$
@@ B.8 Kolmogorov-Arnold neural networks and approximate inverse sampling
$C$ $D$ residual symbolic expression
neg. log $0.001,0.1]^{2} [$$1000,-10]^{2}$ 1.51e-15 $\log(-10.16\,x_{1})-\log(-6.76\,x_{2})+2.23$
Table B14: DLT with KANs: direct sampling, 2-dimensional training regions, residuals and obtained symbolic expressions
@@
${\bf\Phi}_{L}\circ\dots\circ{\bf\Phi}_{1}\colon\mathbb{R}^{d}\to\mathbb{R}^{n}$ (B1)
@@
It is natural to ask whether approximate inverse sampling improves the performance of KANs in determining the symbolic expression of the Legendre transform. We do not have a full answer to this question. In the context of the two-dimensional experiments described in Tables 2, 3, and 4 of the main text, we obtained similar results with and without inverse sampling. A two-dimensional experiment with a skewed sample and direct sampling similar to section 6.4. is described in Table B14.
Table B15: Benchmark results comparing implicit DLT against direct learning with known duals
Function $d$ Model $L^{2}$ Error Time (s)
Impl. Dir. Impl. Dir.
Quadratic 50 MLP 3.55e-03 3.59e-03 87.5 86.6
MLP_ICNN 6.46e+02 6.46e+02 87.3 76.6
ResNet 2.92e-03 3.25e-03 123.8 119.7
ICNN 7.18e-02 6.66e-02 97.4 97.8
100 MLP 2.77e-02 2.46e-02 95.0 95.3
MLP_ICNN 2.54e+03 2.59e+03 96.8 87.4
ResNet 1.88e-02 1.91e-02 136.3 132.0
ICNN 9.59e-02 9.06e-02 105.7 105.5
200 MLP 1.43e-01 1.50e-01 113.5 112.7
MLP_ICNN 1.01e+04 1.01e+04 97.6 113.7
ResNet 9.53e-02 9.40e-02 163.1 164.4
ICNN 4.01e-01 3.73e-01 124.1 125.8
Neg. Log 50 MLP 3.81e-01 2.81e-01 97.5 97.8
MLP_ICNN 2.03e+01 2.03e+01 67.3 70.4
ResNet 1.05e-01 1.23e-01 133.5 130.7
ICNN 2.03e+01 2.06e+01 73.5 61.9
100 MLP 8.56e+$0$ 6.67e+$0$ 106.2 106.1
MLP_ICNN 4.00e+01 3.99e+01 73.0 66.2
ResNet 1.38e+$0$ 1.93e+$0$ 141.5 142.3
ICNN 3.99e+01 3.98e+01 80.3 71.1
200 MLP 3.29e+01 2.68e+01 124.7 125.1
MLP_ICNN 8.21e+01 8.20e+01 68.9 78.7
ResNet 2.92e+01 2.99e+01 168.7 174.1
ICNN 1.70e+01 1.68e+01 131.2 137.0
Neg. Entropy 50 MLP 2.71e-01 3.91e-01 99.6 100.8
MLP_ICNN 7.83e+01 7.87e+01 97.5 104.6
ResNet 6.93e-02 7.27e-02 132.8 135.0
ICNN 9.39e+01 9.39e+01 59.8 69.1
100 MLP 1.77e+01 1.65e+01 106.5 108.7
MLP_ICNN 1.45e+02 1.44e+02 103.5 111.2
ResNet 1.53e+$0$ 1.32e+$0$ 142.0 144.2
ICNN 1.93e+02 1.93e+02 62.9 79.9
200 MLP 4.92e+01 5.07e+01 124.2 127.3
MLP_ICNN 2.87e+02 2.87e+02 125.3 133.3
ResNet 5.08e+01 4.02e+01 94.2 175.4
ICNN 9.43e+01 9.30e+01 131.9 139.2

### v2

@@ B.3.2 Benchmark setup
Table B8: Comparison of DLT with direct learning for known Legendre transforms
Function $d$ Model $L^{2}$ Error Time (s)
DLT Dir DLT Dir
Quadratic 50 MLP 3.55e-03 3.59e-03 87.5 86.6
MLP-ICNN 6.46e+02 6.46e+02 87.3 76.6
ResNet 2.92e-03 3.25e-03 123.8 119.7
ICNN 7.18e-02 6.66e-02 97.4 97.8
100 MLP 2.77e-02 2.46e-02 95.0 95.3
MLP-ICNN 2.54e+03 2.59e+03 96.8 87.4
ResNet 1.88e-02 1.91e-02 136.3 132.0
ICNN 9.59e-02 9.06e-02 105.7 105.5
200 MLP 1.43e-01 1.50e-01 113.5 112.7
MLP-ICNN 1.01e+04 1.01e+04 97.6 113.7
ResNet 9.53e-02 9.40e-02 163.1 164.4
ICNN 4.01e-01 3.73e-01 124.1 125.8
Neg. Log 50 MLP 3.81e-01 2.81e-01 97.5 97.8
MLP-ICNN 2.03e+01 2.03e+01 67.3 70.4
ResNet 1.05e-01 1.23e-01 133.5 130.7
ICNN 2.03e+01 2.06e+01 73.5 61.9
100 MLP 8.56e+$0$ 6.67e+$0$ 106.2 106.1
MLP-ICNN 4.00e+01 3.99e+01 73.0 66.2
ResNet 1.38e+$0$ 1.93e+$0$ 141.5 142.3
ICNN 3.99e+01 3.98e+01 80.3 71.1
200 MLP 3.29e+01 2.68e+01 124.7 125.1
MLP-ICNN 8.21e+01 8.20e+01 68.9 78.7
ResNet 2.92e+01 2.99e+01 168.7 174.1
ICNN 1.70e+01 1.68e+01 131.2 137.0
Neg. Entropy 50 MLP 2.71e-01 3.91e-01 99.6 100.8
MLP-ICNN 7.83e+01 7.87e+01 97.5 104.6
ResNet 6.93e-02 7.27e-02 132.8 135.0
ICNN 9.39e+01 9.39e+01 59.8 69.1
100 MLP 1.77e+01 1.65e+01 106.5 108.7
MLP-ICNN 1.45e+02 1.44e+02 103.5 111.2
ResNet 1.53e+$0$ 1.32e+$0$ 142.0 144.2
ICNN 1.93e+02 1.93e+02 62.9 79.9
200 MLP 4.92e+01 5.07e+01 124.2 127.3
MLP-ICNN 2.87e+02 2.87e+02 125.3 133.3
ResNet 5.08e+01 4.02e+01 94.2 175.4
ICNN 9.43e+01 9.30e+01 131.9 139.2
@@ B.4 Experiments with non-separable functions
Pretrained ICNN: We first train a 2-layer ICNN (128 hidden units per layer) for 50,000 iterations to approximate a target quadratic function, then freeze it and use it as the convex function $f$. This represents a realistic scenario where the target function itself is a learned model from data.
@@
(pretrained to quadratic) ($\approx[-3,3]^{d}$ )
@@
Table B10: Performance of DLT with ResNet approximators for non-separable functions. Absolute RMSE, relative RMSE (normalized by function scale) and training times are reported as mean $\pm$ standard deviation over 5 independent trials.
@@ Results.
The results demonstrate that DLT successfully learns all four non-separable function classes, achieving low RMSE errors across different dimensions. The quadratic SPD function shows the best performance with errors decreasing as dimension increases (from $1.15\times 10^{-2}$ at $d=10$ to $2.01\times 10^{-3}$ at $d=50$), demonstrating DLT’s effectiveness in learning structured covariance patterns. The exponential-minus-linear function presents greater challenges due to its unbounded gradient domain, yet DLT maintains reasonable accuracy. The pretrained ICNN case validates DLT’s ability to approximate learned neural network functions.
@@
B.5 Alternative NN approaches to computing Legendre transforms
DLT approximates the convex conjugate $f^{*}$ on $D=\nabla f(C)$ by leveraging the exact Fenchel–Young identity at gradient-mapped points. The key insight is that for any $x\in C$ with $y=\nabla f(x)$, the identity $f^{*}(y)=\langle x,y\rangle-f(x)$ holds exactly.
@@ B.5 Alternative NN approaches to computing Legendre transforms
DLT provides several critical advantages over the Proxy method:
@@
    Efficient sampling: DLT uses the exact gradient mapping $\nabla f$ for sampling, which is computationally efficient and exact. The Proxy method requires every new sample to pass through the learned inverse network $h_{\vartheta}$, adding computational overhead and potential approximation errors at inference time.
@@
    Training overhead: While both methods may use an inverse mapping for sampling control, the Proxy method critically depends on the quality of this mapping for its training targets, requiring more careful and time-consuming pretraining.
Table B11: DLT with approximate-inverse sampling vs. proxy. $h_{\vartheta}$ RMSE measures prelearned inverse quality; $t_{h}$: inverse pretraining time.
@@ B.6 Hamilton–Jacobi equation and time-parameterized inverse gradient sampling
$\begin{cases}\partial_{t}u+H(\nabla_{x}u(x,t))=0,&(x,t)\in Q,\\ u(x,0)=g(x),&x\in\mathbb{R}^{d},\end{cases}$
@@ B.6.1 The Hopf formula and function definitions
The Hopf formula provides an explicit solution to the Hamilton–Jacobi equation:
$u(x,t)=(g^{*}+tH)^{*}(x)=\sup\limits_{p\in\mathbb{R}^{d}}\{\langle x,p\rangle-g^{*}(p)-tH(p)\}$ (B1)
@@
These provide a more challenging test case:
@@
To approximate the HJ solution $u(x,t)$ we use a time-parameterized network $u_{\theta}(x,t)$ with input $(x,t)$.
@@ B.6.2 Time-parameterized inverse gradient sampling
Denote $v(p,t):=g^{*}(p)+t\,H(p)$. Then the Hopf representation can be written as
$u(x,t)=v(\cdot,t)^{*}(x)=\sup_{p\in\mathbb{R}^{d}}\{\langle x,p\rangle-v(p,t)\}.$
Whenever $v(\cdot,t)$ is differentiable (and, e.g., strictly convex so the maximizer is unique), the maximizer $p=p(x,t)$ is characterized by the first-order optimality condition
$x=\nabla_{p}v(p,t)=\nabla g^{*}(p)+t\,\nabla H(p),\qquad\text{hence}\qquad p=\nabla_{x}u(x,t)=(\nabla_{p}v(\cdot,t))^{-1}(x).$
Time-DLT therefore requires samples of pairs $(x,p)$ satisfying this relation. In the case where $H$ and $g$ are quadratic one has $\nabla_{p}v(p,t)=(1+t)p$, so $x=(1+t)p$ (equivalently $p=x/(1+t)$), which allows direct sampling. In non-quadratic cases the inverse map $(\nabla_{p}v(\cdot,t))^{-1}$ is not available in closed form; to generate samples with a prescribed distribution in $x$-space we learn a time-parameterized inverse-gradient network $\Psi_{\phi}(x,t)\approx(\nabla_{p}v(\cdot,t))^{-1}(x)$. Similar to Section 5.3 we train $\Psi_{\phi}$ using $\|\nabla_{p}v(\Psi_{\phi}(x,t),t)-x\|^{2}_{2}$ and $\|\Psi_{\phi}(\nabla_{p}v(p,t),t)-p\|^{2}_{2}$ as loss funcitons, which only require evaluating $\nabla_{p}v=\nabla g^{*}+t\nabla H$. In the implementation, this provides the inverse-gradient sampling step $p\approx\Psi_{\phi}(x,t)$ (and in the quadratic case the explicit formula $p=x/(1+t)$) used to form training triplets $(x,p,t)$.
B.6.3 Time-DLT implementation
Time-DLT directly approximates the Hopf formula through with neural network $u_{\theta}(x,t)$ that learns:
$u_{\theta}(x,t)\approx u(x,t)=\sup_{p}\left\{\langle x,p\rangle-g^{*}(p)-tH(p)\right\}.$
@@ B.6.3 Time-DLT implementation
$\mathcal{L}=\mathbb{E}_{(x,p,t)}\left[\left|u_{\theta}(x,t)-\left(\langle x,p\rangle-g^{*}(p)-tH(p)\right)\right|^{2}\right].$
The sampling of triplets $(x,p,t)$ depends on the function type: analytical formulas for quadratic functions and the trained inverse network for exponential functions. This approach preserves the structure of the Hopf formula while ensuring computational tractability.
@@ B.6.4 Experimental framework
We evaluate both methods across increasing dimensions and function complexities. We sample $(x,t)\sim\mathrm{Unif}\Big([x_{\min},x_{\max}]^{d}\times[0,t_{max}]\Big)$, with $x_{\min}=-2,x_{max}=2,t_{max}=3$ and report in Tables B12 and B13 the resulted RMSE and PDE residuals across different time moments $t\in\{0.1,0.5,1.0,2.0\}$.
@@ B.6.5 Implementation details
To assess accuracy when an explicit solution is unavailable, we compute reference solutions by evaluating for each point $(x,t)$ a reference value by solving the inner optimization over $p$ in the Hopf representation (B1):
@@
  • Higher dimensions ($2<d\leq 5$): PyTorch gradient-based optimization with Adam
@@ B.6.6 Evaluation metrics
    $\varepsilon_{L^{2}}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}|u_{\theta}(x_{i},t)-u_{\text{ref}}(x_{i},t)|^{2}}$
@@
    $\varepsilon_{\text{PDE}}=\sqrt{\mathbb{E}_{(x,t)}\left[|\partial_{t}u_{\theta}+H(\nabla_{x}u_{\theta})|^{2}\right]}$
@@
    $\varepsilon_{\text{IC}}=\sqrt{\mathbb{E}_{x}\left[|u_{\theta}(x,0)-g(x)|^{2}\right]}$
@@
Table B12: Comparison of DGM and Time-DLT on the Hamilton–Jacobi equation with quadratic initial condition and Hamiltonian
@@
$L^{2}$ Error PDE Res. Time (s) $L^{2}$ Error PDE Res. Time (s) Mean $\sigma$
@@
Notes: $L^{2}$ Error represents the mean squared difference between predicted solutions and the analytical solution $u(x,t)=\frac{\sum x_{i}^{2}}{2(1+t)}$. HJ Error measures the residual of the Hamilton–Jacobi equation $|\partial_{t}u+H(\nabla_{x}u)|$.
Table B13: Comparison of DGM and Time-DLT on the Hamilton–Jacobi equation with quadratic initial condition and exponential Hamiltonian
@@
$L^{2}$ Error PDE Residual Time (s) $L^{2}$ Error PDE Residual Time (s) DGM / DLT
@@
Notes: $L^{2}$ Error quantifies the mean squared deviation from the ground truth solution. PDE Residual is the residual of the Hamilton–Jacobi equation $|\partial_{t}u+H(\nabla_{x}u)|$.
Figure B1: Comparison of DGM and Time-DLT at three different times $t=0.1,0.5,1$ for $d=2.$
@@ B.8 Kolmogorov-Arnold neural networks and approximate inverse sampling
${\bf\Phi}_{L}\circ\dots\circ{\bf\Phi}_{1}\colon\mathbb{R}^{d}\to\mathbb{R}^{n}$